# Revisiting the propagation of highly-energetic gamma rays in the Galaxy

### v1

@@ Abstract
1 Introduction
@@
In addition, fluxes of photons with energies larger than $10$ PeV are predicted from ultra-high-energy cosmic-ray interactions with cosmological photon fields, the so-called cosmogenic photons [26, 27, 28], and from primordial relics such as in top-down models [29, 30, 31], thus super-heavy dark matter [32, 33, 34, 35, 36, 37], Z-burst scenarios [38, 39] or topological defects [40, 41]. Nevertheless, many of these models have been strongly constrained already [42, 43, 44, 45]. In the highest energy bands, the upper limits on the gamma-ray diffuse flux above $\sim 10^{17}\;\text{eV}$ from the Pierre Auger Observatory have been employed [46].
@@
Apart from production mechanisms, gamma rays propagate over astrophysical distances before eventually reaching our observatories. During their travel, they might interact with background photons permeating astrophysical spaces, producing electron-positron pairs, to which from now on we refer simply as electrons. Such electrons are then deflected by astrophysical magnetic fields and, in turn, they might scatter through inverse Compton (IC) background photons up to several TeVs. These processes trigger electromagnetic (EM) cascades over astrophysical distances. In this picture, prompt gamma-ray fluxes could be distorted during their travel towards detectors on Earth. Specifically, their energy spectra could present absorption features and arrival direction maps with spreading counts due to “deflected” gamma rays. The latter might lead to the formation of haloes, i.e. extended gamma-ray counts surrounding the emission cores. Moreover, in this picture, the deflection of electrons in the EM cascade produce delays in the arrival times of gamma rays. These effects have been largely studied in the context of gamma rays propagating over intergalactic lengths [47], leading to constraints on intergalactic magnetic field properties [48, 49, 50, 51]. Propagation effects are fundamental not only in the understanding of source production mechanisms, but also for peculiar indirect searches for axion-like particles [52, 53, 54, 55, 56, 57], dark matter candidates. The goal of this paper is to explore this scenario for gamma rays travelling over galactic length scales.
@@
To achieve our goals, it is crucial to characterise the galactic environment. Besides the cosmic microwave background (CMB) and cosmic radio background (CRB), galactic space is permeated by the interstellar radiation field (ISRF). In intergalactic space, the equivalent of the ISRF is the extragalactic background light (EBL) [58, 59, 60, 61, 62, 63, 64]. Previous works investigated gamma-ray galactic absorption due to pair production on ISRF and CMB by computing the optical depth [65, 66, 67, 68, 69]. They found that gamma-ray fluxes from sources located in the galactic center start to be attenuated at energies above $\sim~{}10$ TeV. Moreover, our Galaxy has a peculiar magnetic field structure, product of thermal electrons pervading interstellar space, large-scale plasma bubbles and small-scale turbulent features due to outflows, then enhanced by hydrodynamic turbulences. Galactic magnetic field (GMF) models are few and carry uncertainties, given the complexity and variety of the measurements involved [70, 71, 72].
@@
This work aims to investigate theoretical scenarios of gamma rays, with energy between 100 GeV and 100 PeV, propagating in our Galaxy, using dedicated Monte Carlo simulation tools. It is organised as follows: in Sec. 2 the simulation settings are described together with gamma-ray propagation theory and galactic environmental properties. In Sec. 3, the observables at Earth position from the simulated gamma-ray sources are examined. Chiefly, spectral and spatial observables are related to the specific source position in the Galaxy, because of background photons and magnetic field spatial models. Sec. 4 outlines the conclusions of this work and open questions to address in future follow-up studies.
2 Simulation of galactic gamma-ray propagation
@@
The Boris push algorithm [75] is employed to solve the equations of motion of the charged particles moving in magnetic field regions. The minimum step size is set to a fraction of the smallest estimated Larmor radius of the propagating (charged) particles. It is computed as $R_{L}=E_{e}/(ecB)$, where $E_{e}$ is the electron energy, $B$ the magnetic field intensity, and $e$ and $c$ are the elementary electric charge and the speed of light, respectively. Secondary electrons have, on average, an energy $E_{e}=E_{\gamma_{0}}/2$, where $E_{\gamma_{0}}$ is the primary gamma-ray energy [^1] A parameter estimating the amount of the primary particle energy channeled to the interaction products is the inelasticity [76].. This choice permits deflections to be accurately computed. Within each simulation step, interaction probabilities are randomly drawn via the Monte Carlo algorithms implemented in CRPropa, which could eventually generate the interaction products.
2.1 Interactions and background photon fields
@@
$\lambda^{-1}_{\text{traj}}(E,z)=\frac{\sum_{i=0}^{N}\lambda^{-1}_{i}(E,z)\cdot l_{i}}{\sum_{i=0}^{N}l_{i}}\,,$ (2.2)
where $\lambda^{-1}_{i}(E,z)$ is the inverse mean free path from Eq. 2.1 computed for the $i$-th photon density, $n_{i}(\epsilon,z)$. Then, $l_{i}$ is the distance covered by the particle in the $i$-th domain.
@@
Figure 1: Pair production inverse mean free path from Eq. 2.1 for CRB (purple dashed-dotted line) from Ref. [92], CMB (solid blue line) and EBL (red dotted line) from Ref. [61]. The ones for the ISRF (F98 model) are computed at three different positions: in the nearby of the galactic center, close to Earth position and $\sim 5\;\text{kpc}$ out of the galactic plane. Dashed-dotted green line is the quantity computed from Eq. 2.2 for the line of sight towards the $S_{>}$ source $\sim 25\;\text{kpc}$ far in Fig. 3 (red ISRF domains). The orange dashed line refers to the line of sight to $S_{1}$ source (blue domains).
In the inset, the violet dashed line is the combined inverse mean free path from CMB and ISRF, the latter in the vicinities of the galactic center.
@@
In the current CRPropa version, position-dependent radiation fields are not implemented. For this reason, to approximate the ISRF model, lines of sight to the sources are divided into different ISRF density regions, i.e. the red/blue boxes in the simulation setup shown in Fig. 3, employing the RestrictToRegion module in CRPropa. In each domain, the ISRF reference density is taken as the closest one to the lines of sight. Eq. 2.2 is adopted to compute the total inverse mean free path for pair production along the line of sight in Fig. 1.
@@
2.2 Galactic magnetic field model
@@
The interplay between regular and turbulent structures might be relevant in interpreting the morphology of the arrival direction of the “deflected" gamma rays. Indeed, observed gamma-ray count maps are expected to be shaped according to the specific magnetic field configuration traversed by electrons in EM cascades. In case of no misalignment between the observer line of sight and the emission direction of the source, if $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}\gg 1$, the count map is expected to show a homogeneous halo around the core of the emission. This fact is due to the procedure of stacking many magnetic field realisations, averaging out random component contributions. In the opposite case or, at least, if $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}\sim 1$, the observed counts map is expected to show asymmetric features due to the effects of the regular component on the charged particles in the EM cascade, in a similar fashion to what happens for extragalactic sources [101]. In App. A magnetic field structures and components are described in more details.
2.3 Spherical observer & gamma-ray sources
@@
In Fig. 1, the weighted inverse mean free paths from Eq. 2.2 are computed for the lines of sight related to the farthest source among $S_{>}$ set, $\sim 25\;\text{kpc}$ distant, and the source $S_{1}$. The former, in particular, is few times smaller than the inverse radius of our Galaxy at around 100 TeV.
@@
3 Results and discussion
This section reports the main results from the simulation of gamma rays injected by each of the ten sources defined above. It is divided into two parts. Sec. 3.1 deals with the energy spectra, whereas Sec. 3.2 concerns gamma rays arrival direction.
3.1 Energy spectra
@@
The two main spectral features are better understood by looking at Fig. 1: the ISRF inverse mean free path for pair production is characterised by a two-peaks profile. The local maximum inverse mean free path at $\sim 1\;\text{TeV}$ is compatible with the first absorption in the energy spectra at $\sim 10\;\text{TeV}$. The absolute maximum in the ISRF inverse mean free path at $\sim 100\;\text{TeV}$ has to be combined with the CMB one, placed at $\sim 10\;\text{PeV}$. Such joint action brings about the maximum absorption at $\sim 2\;\text{PeV}$.
@@
In the case without ISRF, thus including the EBL from Ref. [61], the absorption starts only at 1 PeV due to the interaction of gamma rays with CMB photons (dashed-dotted red line in Fig. 4). The minimum in the spectrum, or maximum absorption, is at $\sim 10\;\text{PeV}$.
@@
Figure 4: Energy spectra for the 10 sources of Fig. 3, both for an injected energy spectrum described by a power law, Eq. 3.1, in (a) and by an exponential cut off, Eq. 3.2, in (b). The spectral energy distributions for the two sets, $S_{>}$ (continuous lines) and $S_{<}$ (dashed), overlap at a fixed distance. The observed energy spectra from $S_{1}$ and $S_{2}$ sources are represented by dotted lines, respectively cyan and sea green. As a reference, the injected energy spectra are given by the continuous dark magenta lines. The energy spectrum for a simulation with the EBL replacing the ISRF is also shown (dashed-dotted red line).
@@
The ratios between the total observed energy and the injected one in dependence to the source-observer distance are shown in Fig. 5, for different emission models. Data points are computed for the case of the power law with spectral index 2. The blue line is the fit referring to $S_{>}$ sources (blue points). A similar result is obtained for $S_{<}$ sources (orange stars), consistently with the energy spectra in Fig. 4. This reference power-law model is compared with larger and smaller spectral indexes and the case with a cut off at 100 TeV, as in Fig. 4(b). Table 1 provides the absorption ratio lengths of the four emission models. The softer the injection model, the smaller the amount of absorbed and/or dispersed energy is. This trend is due to the smaller amount of higher energy photons for larger spectral indexes, i.e. the most affected by absorption and “deflection".
@@
Table 1: In the last column, the slopes of the linear fits in Fig. 5. It reflects the dependence of the distance on the ratio between observed and injected energy. Each row corresponds to a different injection model – PL stands for power-law, in Eq. 3.1, while PLEC for power-law exponentially cut off, Eq. 3.2.
@@ 3.2 Arrival directions
Spatial count maps for the other two sources, $S_{1}$ and $S_{2}$, are respectively shown in Fig. 15 and Fig. 15. The two shapes are strongly irregular both in orientation and in counts distribution around the point-like source at the center. As pointed out in Sec. 2.2, these effects are attributed to the peculiar structure of the magnetic field along the line of sight. In particular, we recall that Fig. 8 and 9 in App. A show, respectively, the magnetic field structures along the line of sight of the $S_{>}$ source $\sim 6\;\text{kpc}$ far from Earth, and the one of the $S_{1}$ source. If the latter line of sight crosses only one regular magnetic field region, the former one is greatly dominated by the turbulent component. In fact, the mean ratio along the line of sight, i.e. $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}$, related to the $S_{1}$ source is about 2.3. In the scenario for the $S_{>}$ source, $\sim 6\;\text{kpc}$ distant, the ratio is $\sim 9.5$. For all the $S_{\lessgtr}$ sources, the line of sight magnetic field structure is extremely turbulent-dominated as well.
@@ 4 Conclusion and prospects
Depending on the specific source position in the Galaxy, gamma-ray fluxes detected at Earth are found to be absorbed and “deflected". The energy spectra shown in Fig. 4 are evidence of absorptions caused by the combination of two photon backgrounds: the CMB and the ISRF. It starts to be relevant at $\sim 10\;\text{TeV}$. In the absence of the ISRF, energy spectra show an absorption at $\sim 1\;\text{PeV}$ due to the CMB. The power-law energy spectrum profiles are in a qualitative agreement with the survival probabilities computed for gamma-ray absorption in Refs. [65, 66, 67, 68, 69], in which ISRF and CMB photons are taken into account. The novelty of our simulations of galactic gamma-ray sources lies in the investigation of EM cascades initiated by such highly energetic fluxes.
@@
The discussion on the arrival direction maps in Sec. 3.2 suggests the complicated detectability of spatially extended haloes around point-like sources due to propagation effects alone. The count maps of this work reveal the high contrast – at least four or five orders of magnitude – between the centered point-like sources and the deflected counts around. Yet, two interesting considerations are drawn. The first one concerns the role of the GMF and how it affects the arrival directions of gamma rays. In the case of galactic TeV-extended sources, e.g. [102, 103], propagation effects might be relevant in enhancing the extension, slightly deforming the shape itself. The latter feature is due to the complex structure of the GMF between source and observer. How much these effects impact current and future gamma-ray observations is left for a future work. The second consideration is about the influence of “deflected” gamma rays in our observation. We enquire whether the angular spreading of gamma rays due to propagation effects contribute to the galactic diffuse gamma-ray background [104, 105, 106, 107, 108, 109, 110], considering the so far observed multi-TeV sources of our Galaxy, viz. the ones in LHAASO and HAWC catalogs [111, 112, 113]. In a broader multi-messenger view, a potential impact on the galactic diffuse gamma-ray background might improve the diffuse neutrino background expectations from gamma rays [114]. Another interesting concern regards the role of the electrons generated in the EM cascades, since their plausible contribution to the diffuse electron background detected at Earth. It might cause anisotropies related to the GMF structure [115, 116, 117].
@@
Overall, the significant point would be to compare what learnt from simulations, in terms of gamma-ray absorption and “deflection", with current and future gamma-ray experiments performance, to properly infer the impact of propagation effects on observed galactic gamma-ray fluxes. This will result in improved characterisations of PeVatrons and galactic center intrinsic emission models, as well as in more robust constraints from dark matter searches through gamma rays.
Acknowledgments

### v4

@@ Abstract
gamma-ray propagation, galactic environment: magnetic field and background photons
preprint: APS/123-QED
I Introduction
@@
In addition, fluxes of photons with energies larger than $10$ PeV are predicted from ultra-high-energy cosmic-ray interactions with cosmological photon fields, the so-called cosmogenic photons [26, 27, 28], and from primordial relics such as in top-down models [29, 30, 31], thus super-heavy dark matter [32, 33, 34, 35, 36, 37, 38], Z-burst scenarios [39, 40] or topological defects [41, 42]. Nevertheless, many of these models have been strongly constrained already [43, 44, 45, 46]. In the highest energy bands, the upper limits on the gamma-ray diffuse flux above $\sim 10^{17}\;\text{eV}$ from the Pierre Auger Observatory have been employed [47].
@@
Apart from production mechanisms, gamma rays propagate over astrophysical distances before eventually reaching our observatories. During their travel, they might interact with background photons permeating astrophysical spaces, producing electron-positron pairs, to which from now on we refer simply as electrons. Such electrons are then deflected by astrophysical magnetic fields and, in turn, they might scatter through inverse Compton (IC) background photons up to several TeVs. These processes trigger electromagnetic (EM) cascades over astrophysical distances [^1] Throughout the text, even with a low number of occurring interactions we refer to the phenomena as EM cascade.. In this picture, prompt gamma-ray fluxes could be distorted during their travel towards detectors on Earth. Specifically, their energy spectra could present absorption features and arrival direction maps with spreading counts due to “deflected” gamma rays. The latter might lead to the formation of haloes, i.e. extended gamma-ray counts surrounding the emission cores. Moreover, in this picture, the deflection of electrons in the EM cascade produce delays in the arrival times of gamma rays. These effects have been largely studied in the context of gamma rays propagating over intergalactic lengths [48], leading to constraints on intergalactic magnetic field properties [49, 50, 51, 52]. Propagation effects are fundamental not only in the understanding of source production mechanisms, but also for peculiar indirect searches for axion-like particles [53, 54, 55, 56, 57, 58], dark matter candidates. The goal of this paper is to explore this scenario for gamma rays travelling over galactic length scales.
@@
To achieve our goals, it is crucial to characterise the galactic environment. Besides the cosmic microwave background (CMB) and cosmic radio background (CRB), galactic space is permeated by the interstellar radiation field (ISRF). In intergalactic space, the equivalent of the ISRF is the extragalactic background light (EBL) [59, 60, 61, 62, 63, 64, 65]. Previous works investigated gamma-ray galactic absorption due to pair production on ISRF and CMB by computing the optical depth [66, 67, 68, 69, 70, 71]. They found that gamma-ray fluxes from sources located in the galactic center start to be attenuated at energies above $\sim~{}10$ TeV. Moreover, our Galaxy has a peculiar magnetic field structure, product of thermal electrons pervading interstellar space, large-scale plasma bubbles and small-scale turbulent features due to outflows, then enhanced by hydrodynamic turbulences. Galactic magnetic field (GMF) models are few and carry uncertainties, given the complexity and variety of the measurements involved [72, 73, 74, 75].
@@
This work aims to investigate theoretical scenarios of gamma rays, with energy between 100 GeV and 100 PeV, propagating in our Galaxy, using dedicated Monte Carlo (MC) simulation tools. It is organised as follows: in Sec. II the simulation settings are described together with gamma-ray propagation theory and galactic environmental properties. In Sec. III, the observables at Earth position from the simulated gamma-ray sources are examined. Chiefly, spectral and spatial observables are related to the specific source position in the Galaxy, because of background photons and magnetic field spatial models. Sec. IV outlines the conclusions of this work and open questions to address in future follow-up studies.
II Simulation of galactic gamma-ray propagation
@@
The Boris push algorithm [78] is employed to solve the equations of motion of the charged particles moving in magnetic field regions. The minimum step size is set to a fraction of the smallest estimated Larmor radius of the propagating (charged) particles. It is computed as $R_{L}=E_{e}/(ecB)$, where $E_{e}$ is the electron energy, $B$ the magnetic field intensity, and $e$ and $c$ are the elementary electric charge and the speed of light, respectively. This choice permits deflections to be accurately computed. Secondary electrons have, on average, an energy $E_{e}=E_{\gamma_{0}}/2$, where $E_{\gamma_{0}}$ is the primary gamma-ray energy. The parameter estimating the amount of the primary particle energy channeled to the interaction products is the inelasticity [79], that can be very large for EM processes. Within each simulation step, interaction probabilities are randomly drawn via the MC algorithms implemented in CRPropa, which could eventually generate the interaction products.
II.1 Interactions and background photon fields
@@
$\lambda^{-1}_{\text{traj}}(E,z)=\frac{\sum_{i=0}^{N}\lambda^{-1}_{i}(E,z)\cdot l_{i}}{\sum_{i=0}^{N}l_{i}}\,,$ (2)
where $\lambda^{-1}_{i}(E,z)$ is the inverse mean free path from Eq. 1 computed for the $i$-th photon density, $n_{i}(\epsilon,z)$. Then, $l_{i}$ is the distance covered by the particle in the $i$-th domain.
@@
Figure 1: Pair production inverse mean free path from Eq. 1 for CRB (purple dashed-dotted line) from Ref. [95], CMB (solid blue line) and EBL (red dotted line) from Ref. [62]. The ones for the ISRF (F98 model) are computed at three different positions: in the nearby of the galactic center, close to Earth position and $\sim 5\;\text{kpc}$ out of the galactic plane. The dashed-dotted green line is the quantity computed from Eq. 2 for the line of sight towards the $S_{>}$ source $\sim 25\;\text{kpc}$ far in Fig. 3 (red ISRF domains). The orange dashed line refers to the line of sight to $S_{1}$ source (blue domains). In the inset, the violet dashed line is the combined inverse mean free path from CMB and ISRF, the latter in the vicinities of the galactic center.
@@
In the current CRPropa version, position-dependent radiation fields are not implemented. For this reason, to approximate the ISRF model, lines of sight to the sources are divided into different ISRF density regions, i.e. the red/blue boxes in the simulation setup shown in Fig. 3, employing the RestrictToRegion module in CRPropa. In each domain, the ISRF reference density is taken as the closest one to the lines of sight. Eq. 2 is adopted to compute the total inverse mean free path for pair production along the line of sight in Fig. 1.
@@
II.2 Galactic magnetic field model
@@
The interplay between regular and turbulent structures might be relevant in interpreting the morphology of the arrival direction of the “deflected” gamma rays. Indeed, observed gamma-ray count maps are expected to be shaped according to the specific magnetic field configuration traversed by electrons in EM cascades. In case of no misalignment between the observer line of sight and the emission direction of the source, if $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}\gg 1$, the count map is expected to show a homogeneous halo around the core of the emission. This fact is due to the procedure of stacking many magnetic field realisations, averaging out random component contributions. In the opposite case or, at least, if $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}\sim 1$, the observed counts map is expected to show asymmetric features due to the effects of the regular component on the charged particles in the EM cascade, in a similar fashion to what happens for extragalactic sources [104]. In App. A magnetic field structures and components are described in more details.
II.3 Spherical observer & gamma-ray sources
@@
In Fig. 1, the weighted inverse mean free paths from Eq. 2 are computed for the lines of sight related to the farthest source among $S_{>}$ set, $\sim 25\;\text{kpc}$ distant, and the source $S_{1}$. The former, in particular, is few times smaller than the inverse radius of our Galaxy at around 100 TeV.
@@
III Results and discussion
This section reports the main results from the simulation of gamma rays injected by each of the ten sources defined above. It is divided into two parts. Sec. III.1 deals with the energy spectra, whereas Sec. III.2 concerns gamma rays arrival direction.
III.1 Energy spectra
@@
The two main spectral features are better understood by looking at Fig. 1: the ISRF inverse mean free path for pair production is characterised by a two-peaks profile. The local maximum inverse mean free path at $\sim 1\;\text{TeV}$ is compatible with the first absorption in the energy spectra at $\sim 10\;\text{TeV}$. The absolute maximum in the ISRF inverse mean free path at $\sim 100\;\text{TeV}$ has to be combined with the CMB one, placed at $\sim 10\;\text{PeV}$. The maximum absorption is located at $\sim 2\;\text{PeV}$.
@@
In the case without ISRF, thus including the EBL from Ref. [62], the absorption starts only at 1 PeV due to the interaction of gamma rays with CMB photons (dashed-dotted red line in Fig. 6). The minimum in the spectrum, or maximum absorption, is at $\sim 2\;\text{PeV}$ as well.
Figure 4: *
@@
Figure 6: Energy spectra for the 10 sources of Fig. 3, both for an injected energy spectrum described by a power law, Eq. 3, in (a) and by an exponential cut off, Eq. 4, in (b). The spectral energy distributions for the two sets, $S_{>}$ (continuous lines) and $S_{<}$ (dashed), overlap at a fixed distance. The observed energy spectra from $S_{1}$ and $S_{2}$ sources are represented by dotted lines, respectively cyan and sea green. As a reference, the injected energy spectra are given by the continuous dark magenta lines. The energy spectrum for a simulation with the EBL replacing the ISRF is also shown (dashed-dotted red line). In this latter scenario, we simulate gamma rays only up to 10 PeV.
@@
The ratios between the total observed energy and the injected one in dependence to the source-observer distance are shown in Fig. 7, for different emission models. Data points are computed for the case of the power law with spectral index 2. The blue line is the fit referring to $S_{>}$ sources (blue points). A similar result is obtained for $S_{<}$ sources (orange stars), consistently with the energy spectra in Fig. 6. This reference power-law model is compared with larger and smaller spectral indexes and the case with a cut off at 100 TeV, as in Fig. 6(b). Table 1 provides the absorption ratio lengths of the four emission models. The softer the injection model, the smaller the amount of absorbed and/or dispersed energy is. This trend is due to the smaller amount of higher energy photons for larger spectral indexes, i.e. the most affected by absorption and “deflection”.
@@
Table 1: In the last column, the slopes of the linear fits in Fig. 7. It reflects the dependence of the distance on the ratio between observed and injected energy. Each row corresponds to a different injection model – PL stands for power-law, in Eq. 3, while PLEC for power-law exponentially cut off, Eq. 4.
@@ III.2 Arrival directions
Spatial count maps for the other two sources, $S_{1}$ and $S_{2}$, are respectively shown in Fig. 29 and Fig. 29. The two shapes are strongly irregular both in orientation and in counts distribution around the point-like source at the center. As pointed out in Sec. II.2, these effects are attributed to the peculiar structure of the magnetic field along the line of sight. In particular, we recall that Fig. 14 and 15 in App. A show, respectively, the magnetic field structures along the line of sight of the $S_{>}$ source $\sim 6\;\text{kpc}$ far from Earth, and the one of the $S_{1}$ source. If the latter line of sight crosses only one regular magnetic field region, the former one is greatly dominated by the turbulent component. In fact, the mean ratio along the line of sight, i.e. $\langle B_{\text{turb}}/B_{\text{reg}}\rangle_{\text{los}}$, related to the $S_{1}$ source is about 2.3. In the scenario for the $S_{>}$ source, $\sim 6\;\text{kpc}$ distant, the ratio is $\sim 9.5$. For all the $S_{\lessgtr}$ sources, the line of sight magnetic field structure is extremely turbulent-dominated as well.
@@ IV Conclusion and prospects
Depending on the specific source position in the Galaxy, gamma-ray fluxes detected at Earth are found to be absorbed and “deflected”. The energy spectra shown in Fig. 6 are evidence of absorptions caused by the combination of two photon backgrounds: the CMB and the ISRF. It starts to be relevant at $\sim 10\;\text{TeV}$. In the absence of the ISRF, energy spectra show an absorption at $\sim 1\;\text{PeV}$ due only to the CMB. The power-law energy spectrum profiles are in a qualitative agreement with the survival probabilities computed for gamma-ray absorption in Refs. [66, 67, 69, 70, 71], in which ISRF and CMB photons are taken into account. The novelty of our simulations of galactic gamma-ray sources lies in the spectral and spatial investigation of EM cascades initiated by such highly energetic fluxes, through MC simulation tools.
@@
The discussion on the arrival direction maps in Sec. III.2 suggests the complicated detectability of spatially extended haloes around point-like sources due to propagation effects alone – even though they do depend on the intrinsic gamma-ray spectrum of the source. The count maps showed in this work reveal the high contrast – at least four or five orders of magnitude – between the centered point-like sources and the deflected counts around. Yet, two interesting considerations are drawn. The first one concerns the role of the GMF and how it affects the arrival directions of gamma rays. In the case of galactic TeV-extended sources, e.g. [105, 106], propagation effects might be relevant in enhancing the extension, slightly deforming the shape itself. The latter feature is due to the complex structure of the GMF between source and observer. How much these effects impact current and future gamma-ray observations is left for a future work. The second consideration is about the influence of “deflected” gamma rays in our observation. We enquire whether the angular spreading of gamma rays due to propagation effects contribute to the galactic diffuse gamma-ray background [107, 108, 109, 110, 111, 112, 113], considering the so far observed multi-TeV sources of our Galaxy, viz. the ones in LHAASO and HAWC catalogs [114, 115, 116]. In a broader multi-messenger view, a potential impact on the galactic diffuse gamma-ray background might improve the diffuse neutrino background expectations from gamma rays [117]. Another interesting concern regards the role of the electrons generated in the EM cascades, since their plausible contribution to the diffuse electron background detected at Earth. It might cause anisotropies related to the GMF structure [118, 119, 120].
@@
Overall, the significant point would be to compare what learnt from simulations, in terms of gamma-ray absorption and “deflection”, with current and future gamma-ray experiments performance, to properly infer the impact of propagation effects on observed galactic gamma-ray fluxes. This will result in improved characterisations of PeVatrons and galactic center intrinsic emission models, as well as in more robust constraints from dark matter searches through gamma rays.
Acknowledgements.